OpenStudy (anonymous):

What is the slope of the line that passes through the points (1, 3) and (5, -2)

OpenStudy (shamil98):

use the slope formula \[\huge m = \frac{ y_2 - y_1 }{ x_2 - x_1 }\]

OpenStudy (anonymous):

Okay... Help me step by step?

OpenStudy (anonymous):

Slope formula \[\frac{ y2 - y1 }{ x2 -x1 }\] or \[\frac{ y1 - y2 }{ x1 - x2 }\] The two points (1,3),(5,-2) (x1,y1),(x2,y2) or (1,3),(5,-2) (x2,y2),(x1,y1) They are interchangeable but must correspond correctly when plugged in. \[\frac{ 3 -(-2) }{ 1 - 5 }\] or \[\frac{ -2 - 3 }{ 5 - 1 }\] You get -5/4 regardless of which way you work it

OpenStudy (shamil98):

Plug in the numbers, what do you get? (set it up)

OpenStudy (anonymous):

Oh that's how you figure that out thank you :D both of you